At t= 0, a particle starts from rest at z= 0, y= 0, and moves in the zy plane with an acceleration à = (4.0 i + 3.03)m/s2 Assume t is in seconds.
A) Calculate the x and ycomponent of velocity at 2.0s. B) Determine the speed of the particle at 2.0s. C) Determine the position of the particle at 2.0s.
Answer #1

here,

acceleration , a = ( 4 i + 3 j) m/s^2

at t = 2 s

the velocity , v = u + a * t

v = 0 + ( 4 i + 3 j) * 2

v = (8 i + 6 j ) m/s

the x-component of velocity is 8 m/s

the y-component of velocity is 6 m/s

B)

the speed of particle , |v| = sqrt(8^2 + 6^2)

|v| = 10 m/s

c)

the position of particle , r = v0 * t + 0.5 * a * t^2

r = 0 + 0.5 * ( 4 i + 3 j) * 2^2

r = ( 8 i + 6 j) m